# Data-Driven Discovery of Mobility Periodicity for Understanding Urban Transportation Systems

### v1

@@ Abstract
Uncovering the temporal regularity of human mobility is crucial for discovering urban dynamics and has implications for various decision-making processes and urban system applications. This study formulates the periodicity quantification problem in complex and multidimensional human mobility data as a sparse identification of dominant positive auto-correlations in time series autoregression, allowing one to discover and quantify significant periodic patterns such as weekly periodicity from a data-driven and interpretable machine learning perspective. We apply our framework to real-world human mobility data, including metro passenger flow in Hangzhou, China and ridesharing trips in New York City (NYC) and Chicago, USA, revealing the interpretable weekly periodicity across different spatial locations over past several years. In particular, our analysis of ridesharing data from 2019 to 2024 demonstrates the disruptive impact of the COVID-19 pandemic on mobility regularity and the subsequent recovery trends, highlighting differences in the recovery pattern percentages and speeds between NYC and Chicago. We explore that both NYC and Chicago experienced a remarkable reduction of weekly periodicity in 2020, and the recovery of mobility regularity in NYC is faster than Chicago. The interpretability of sparse autoregression provides insights into the underlying temporal patterns of human mobility, offering a valuable tool for understanding urban systems. Our findings highlight the potential of interpretable machine learning to unlock crucial insights from real-world mobility data.
@@ 1 Introduction
Human mobility in cities often reveals recurring patterns—the predictable ebbs and flows of people moving through urban areas—that reflect the underlying structure of urban life (Gonzalez et al., 2008; Song et al., 2010; Simini et al., 2012). From daily commutes to weekend trips, these temporal regularities in human mobility underpin everything from urban transportation management (Sheffi, 1985) and transit scheduling (Guihaire and Hao, 2008) to predicting the spread of diseases (Balcan et al., 2009; Belik et al., 2011; Castillo-Chavez et al., 2016; Li et al., 2022; Du et al., 2018) and optimizing resource allocation (Katoh et al., 2024). In essence, human mobility regularity is an important concept for identifying periodic trips in urban areas, which often follow daily and weekly cycles. A clear example is daily commuting, where passenger flows typically exhibit morning and afternoon peaks, forming a roughly “M"-shaped pattern in time series data (Chen and Sun, 2022) (see e.g., Fig. 1B). As cities evolve, mobility patterns have become increasingly complex and less predictable, characterized by complicated regulatory policies, diverse transportation modes and behavioral shifts driven by social economic and environmental factors (Midgley, 2009; Machado et al., 2018; Tirachini, 2020; Nouvellet et al., 2021; Li et al., 2022). Thus, understanding the rhythm of human mobility in urban areas is essential for grasping how our societies function and urban systems operate (Gonzalez et al., 2008; Song et al., 2010; Simini et al., 2012; Batty et al., 2012; Jiang et al., 2017; Alessandretti et al., 2020).
@@
However, these approaches typically do not provide an explicit and interpretable metric that directly quantifies periodicity, making them less suitable for downstream modeling or individualized interpretation. In addition, most studies that quantify periodicity or regularity in human mobility data focus on the analysis of individual trajectory data rather than aggregated demand or travel-mode usage data (Prabhala et al., 2014; Goulet-Langlois et al., 2017; Huang et al., 2019; Teixeira et al., 2021; do Couto Teixeira et al., 2021; Tao et al., 2021; Manley et al., 2018; Oliveira et al., 2016). While individual data can provide useful information on predicting mobility flows, interpretable information on aggregated demand data reveals longer term trends that can be more directly related to larger economic, social, and other changes that can help inform the development of policies in the future. Another fact is that mobility flows can be severely disrupted by extreme events and may evolve gradually over time due to changes in behavior, infrastructure, or policy. For example, the COVID-19 pandemic disrupted regular travel routines and led to a noticeable decline in mobility regularity (Li et al., 2022; Chen et al., 2025a). The challenge is how to systematically use periodicity strength to detect, quantify, and interpret such changes across both space and time.
@@
Figure 1: Human mobility data and their time series periodicity. (A) Human mobility data can be structured as algebraic tensors where each entry corresponds to the value of a specific variable $\gamma$ at a particular spatial location $n$ and time step $t$. Here, $N$ represents the number of spatial locations, and $\delta$ refers to the number of variables or time phases. The variable dimension $\gamma$ is flexible and can encode different aspects of mobility, such as inflow and outflow at a particular spatial location. Moreover, the temporal lags such as data from different years can also be incorporated as distinct variables within this tensor framework. (B) Illustration of weekly periodicity in passenger flow time series. This panel displays an example of an hourly passenger flow time series that exhibits a clear weekly periodicity. This periodicity is visually highlighted by annotating the data points according to weekly cycles (e.g., blue and orange circles). The time series reveals distinct mobility patterns between weekdays and weekends, with weekdays showing more pronounced morning and afternoon peak hours. (C) Example time series of passenger flow for comparison. (D) Auto-correlation analysis through scatter plots. These panels present scatter plots that visualize the relationship between trip counts at a given time step $t$ and trip counts at previous time steps $t-k$, in which $k$ represents different time lags (e.g., $k=1,24,168$ correspond to hourly, daily, and weekly cycles, respectively). Points that align closely along the anti-diagonal in these scatter plots indicate a strong similarity between the trip counts at the two time steps, implying positive auto-correlation and potential periodicity. Notably, the scatter plot for a weekly cycle demonstrates a more remarkable similarity compared to the daily and hourly cycles, visually confirming the strong weekly periodicity in the example time series.
@@
Substantial progress has been made in understanding human mobility regularity and patterns from analytical perspectives such as entropy and predictability (Oliveira et al., 2016; Barbosa et al., 2018; Delussu et al., 2023). However, the existing methods are not well-suited to quantifying periodicity across spatiotemporal domains and discovering dynamic patterns that related to human mobility regularity. Recently, interpretable machine learning methods (Murdoch et al., 2019; Rudin et al., 2022; Brunton and Kutz, 2022) offer new opportunities to improve appearance, robustness, and interpretability of data-driven exploration in mobility data. In terms of model development, sparse linear regression (Tibshirani, 1996; Jenatton et al., 2011) can be used to reformulate the problem of sparse identification of nonlinear dynamical systems (Brunton et al., 2016), which laid the foundations of several machine learning frameworks that designed for variable and feature selection in scientific discovery. In that venue, sparsity is one of the most important aspects to reinforce interpretability of machine learning models (Murdoch et al., 2019; Rudin et al., 2022). However, existing sparse methods face critical limitations when applied to large-scale human mobility data. First, classical sparse regression methods are primarily designed for variable selection from univariate and multivariate data and struggle to scale to multidimensional data with spatial and temporal dependencies. Second, mobility data are inherently noisy, irregular, and heterogeneous across spatiotemporal domains, which challenges standard sparse models that lack built-in mechanisms for structured robustness. Third, even when sparsity in enforced, the resulting models do not yield interpretable or comparable metrics for periodicity strength across different spatial locations, variables, and time phases.
@@
To bridge these gaps, we propose a multidimensional sparse autoregression framework that models time series of mobility flows across space and time, explicitly captures dominant auto-correlations, and provides a unified and interpretable measure of periodicity strength. This interpretable machine learning framework builds upon the tensor data structure. Formally, human mobility data are spatially- and temporally-varying, and can be represented as a third-order tensor of non-negative integers $x_{n,\gamma,t}\in\{0,1,2,\cdots\}$ as shown in Fig. 1A, where $n\in[N]$ indexes spatial locations, $\gamma\in[\delta]$ indexes different variables or time phases, and $t\in[T_{\gamma}]$ indexes time steps. This algebraic tensor structure captures the multidimensional nature of mobility systems and aligns with a rich body of work on tensor-based machine learning for spatiotemporal pattern discovery (Kolda and Bader, 2009; Chen et al., 2024). Each individual sequence $x_{n,\gamma,t}$ is a time series, and thus can be seamlessly characterized by using time series models such as autoregression (Hamilton, 2020). As exemplified by Fig. 1B, real-world passenger flow time series demonstrate weekly periodicity by picking data points with a weekly cycle. When these data points are denoted by $x_{t},t\in[T]$, the auto-correlations in Fig. 1D can be formulated as $x_{t}\approx w_{k}x_{t-k}$, measuring self-similarity in mobility over any time lag $k$. Observing Fig. 1D, data points of the example time series are very close to the anti-diagonal curve at a weekly cycle (i.e., $k=168$), implying that weekly periodicity is positively auto-correlated. Moreover, the “closeness" of data points to the anti-diagonal at index $k=168$ in Fig. 1D justifies the strength of weekly periodicity of time series, demonstrating that the time series in Fig. 1C is less periodic at a weekly cycle than Fig. 1B.
@@ 2 Machine Learning Formula
Figure 2: Illustration of time series autoregression and its multidimensional setting on human mobility data. (A) The $d$th-order autoregression on the univariate time series can be formulated as an optimization problem for finding the optimal coefficients. Here, the order $d$ is an integer, determining the number of auto-correlations. (B) Multidimensional time series autoregression. Note that the symbol “$\times$" simply refers to the element-wise product between matrices. (C) The interpretability of time series autoregression can be reinforced by sparse and non-negative coefficients. The upper bound of the number of nonzero entries in the coefficient vector is set as an integer $\tau$, which is formally defined as the sparsity level. In mathematics, the cardinality of index set $\Omega$ (i.e., indices of nonzero coefficients) can be used to reformulate sparsity constraints. As a result, the decision variables in the constructed optimization problem include both coefficients and index set.
Interpretable machine learning provides a way to understand underlying patterns of complicated data and systems (Brunton et al., 2016; Murdoch et al., 2019). This study starts from the core idea of autoregression on time series, namely, the current value of the time series can be predicted as a linear combination of its past values (Hamilton, 2020). The objective is to find a set of coefficients that minimizes the sum of squared autoregressive errors (Fig. 2A) with predefined constraints such as sparsity and non-negativity on the coefficients as shown in Fig. 2C. Sparsity, enforced by the constraint $|\Omega|\leq\tau$, limits the number of nonzero coefficients to a small integer $\tau$. This implies that only a few past time lags significantly contribute to predicting the current value. For instance, Fig. 1D shows that the trip counts at time $t$ are highly consistent with the trip counts at time $t-168$, revealing predictable and periodic patterns.
@@
where $w_{n,\gamma,k},n\in[N],\gamma\in[\delta],k\in[d]$ are the non-negative coefficients of this $d$th-order autoregression formula, allowing one to quantify the positive auto-correlations such as daily and weekly periodicity as the dominant ones in time series. The last constraint is the normalization on the non-negative coefficients, letting the sum of the coefficients be one. Since the optimization problem is equivalent to the $\ell_{0}$-norm induced sparse autoregression, it can be addressed by mixed-integer programming solvers with exact solutions (Bertsimas et al., 2016, 2020, 2024; Tillmann et al., 2024). On such multidimensional tensor time series, we can discover the weekly periodicity of human mobility across different spatial locations and variables (or time phases).
@@ 3 Metro Passenger Flow Periodicity
To demonstrate the effectiveness of our framework in capturing periodicity from real-world mobility data, we begin with a case study of metro passenger flow data in Hangzhou, China. This dataset offers a clean and well-structured setting with strong known periodic patterns, making it an ideal testbed for validating the ability of our sparse autoregression model to recover meaningful temporal cycles—particularly weekly periodicity across different stations and flow directions.
@@
Figure 3: Identification of weekly periodicity in inflow and outflow time series from the Hangzhou metro passenger flow dataset (from January 2 to 25, 2019) with the multidimensional sparse autoregression method (sparsity level $\tau=4$). (A) Daily average inflow trips of 81 metro stations. Note that the trip records of the station #54 are missing. (B) Weekly periodicity of hourly inflow time series achieved by the sparse autoregression with sparsity level $\tau=4$. 64 stations show high periodicity ($\geq$ 0.8). The average periodicity value is 0.844. (C) Weekly periodicity of hourly outflow time series. 53 stations show high periodicity ($\geq$ 0.8). The average periodicity value is 0.829. Since inflow and outflow are two variables in the tensor data, their coefficients are optimized simultaneously with the same index set. (D) Scatter plots of trip counts at time $t$ versus time $t-168$ of 6 selected stations, illustrating varying strengths of weekly periodicity. The time series with a high weekly periodicity value also presents a strong similarity of data points between times $t$ and $t-168$. (E) Hourly inflow time series for stations #14 and #15, shown across three consecutive weeks: January 2–8, January 9–15, and January 16–22. Strong overlap among these weekly segments indicates high periodicity at a weekly cycle.
[…]
@@
Table 1: Strength of weekly periodicity of the Hangzhou metro passenger flow data at the selected end stations of metro lines.
Station # Inflow Outflow Station # Inflow Outflow
23 0.77 0.65 35 0.91 0.76
24 0.46 0.48 36 0.88 0.74
25 0.24 0.54 37 0.85 0.74
26 0.78 0.62 38 0.76 0.75
27 0.93 0.78 39 0.73 0.74
28 0.87 0.69 40 0.83 0.88
29 0.77 0.66 62 0.95 0.82
30 0.81 0.71 63 0.89 0.77
31 0.90 0.70 64 0.89 0.81
32 0.91 0.74 65 0.86 0.80
33 0.90 0.80 66 0.87 0.73
34 0.90 0.76 67 0.93 0.79
@@
Fig. 3D shows the weekly periodicity values for six selected stations, measured as the strength of auto-correlations at a weekly lag (i.e., $k=168$) in the inflow time series. The source emphasizes that time series with high weekly periodicity values also show a strong similarity between data points at times $t$ and $t-168$, indicating positive auto-correlations. The weekly periodicity is visually confirmed by scatter plots and time series overlaps. For instance, station #14 shows a high weekly periodicity of 0.93, and its scatter plot shows a strong alignment along the anti-diagonal (Fig. 3D). In particular, the comparison between stations #14 and #15 in Fig. 3E answers why the strength of time series periodicity is a fair metric to reveal repeating patterns of human mobility. The periodicity value of station #15 is 0.71, which is smaller than station #14. As shown in Fig. 3E, the trip count time series at station #15 does not align perfectly across weeks, indicating weaker weekly periodicity compared to the more consistent pattern observed at station #14. To summarize, on the passenger flow dataset, the comparable metric of weekly periodicity across different stations and inflow/outflow variables to discover periodic mobility patterns in a city-wide scale.

### v2

# Data-Driven Discovery of Mobility Periodicity for Understanding Urban Systems

@@ Abstract
Human mobility regularity is crucial for understanding urban dynamics and informing decision-making processes. This study first quantifies the periodicity in complex human mobility data as a sparse identification of dominant positive auto-correlations in time series autoregression and then discovers periodic patterns. We apply the framework to large-scale metro passenger flow data in Hangzhou, China and multi-modal mobility data in New York City and Chicago, USA, revealing the interpretable weekly periodicity across different spatial locations over past several years. The analysis of ridesharing data from 2019 to 2024 demonstrates the disruptive impact of the pandemic on mobility regularity and the subsequent recovery trends. In 2024, the periodic mobility patterns of ridesharing, taxi, subway, and bikesharing in Manhattan uncover the regularity and variability of these travel modes. Our findings highlight the potential of interpretable machine learning to discover spatiotemporal mobility patterns and offer a valuable tool for understanding urban systems.
@@ 1 Introduction
Human mobility in cities often reveals recurring patterns—the predictable ebbs and flows of people moving through urban areas—that reflect the underlying structure of urban life (Gonzalez et al., 2008; Song et al., 2010; Simini et al., 2012). From daily commutes to weekend trips, these temporal regularities in human mobility underpin everything from urban transportation management (Sheffi, 1985) and transit scheduling (Guihaire and Hao, 2008) to predicting the spread of diseases (Balcan et al., 2009; Belik et al., 2011; Castillo-Chavez et al., 2016; Li et al., 2022; Du et al., 2018) and optimizing resource allocation (Katoh et al., 2024). In essence, human mobility regularity is an important concept for identifying periodic trips in urban areas, which often follow daily and weekly cycles. A clear example is daily commuting, where passenger flows typically exhibit morning and afternoon peaks, forming a roughly “M”-shaped pattern in trip time series data (Chen and Sun, 2022) (see e.g., Fig. 1B). As cities evolve, mobility patterns have become increasingly complex and less predictable, characterized by complicated regulatory policies, diverse transportation modes and behavioral shifts driven by social economic and environmental factors (Midgley, 2009; Machado et al., 2018; Tirachini, 2020; Nouvellet et al., 2021; Li et al., 2022). Thus, understanding the rhythm of human mobility in urban areas is essential for grasping how our societies function and urban systems operate (Gonzalez et al., 2008; Song et al., 2010; Simini et al., 2012; Batty et al., 2012; Jiang et al., 2017; Alessandretti et al., 2020).
@@
However, these approaches typically do not provide an explicit and interpretable metric that directly quantifies periodicity, making them less suitable for downstream modeling or individualized interpretation. In addition, most studies that quantify periodicity or regularity in human mobility data focus on the analysis of individual trajectory data rather than aggregated demand or travel-mode usage data (Prabhala et al., 2014; Goulet-Langlois et al., 2017; Huang et al., 2019; Teixeira et al., 2021; do Couto Teixeira et al., 2021; Tao et al., 2021; Manley et al., 2018; Oliveira et al., 2016). While individual data can provide useful information on predicting mobility flows, interpretable information on aggregated demand data reveals longer term trends that can be more directly related to larger economic, social, and other changes that can help inform the development of policies in the future. In urban systems, human mobility patterns vary with different travel modes such as public transit, subway, ridesharing, and bikesharing. It is also meaningful to fairly quantify mobility regularity of different travel modes and utilize periodicity patterns to allocate resources and inform urban planning. Another fact is that mobility flows can be severely disrupted by extreme events and may evolve gradually over time due to changes in behavior, infrastructure, or policy. For example, the COVID-19 pandemic disrupted regular travel routines and led to a noticeable decline in mobility regularity (Li et al., 2022; Chen et al., 2025a). The challenge is how to systematically use periodicity strength to detect, quantify, and interpret such changes across both space and time.
@@
Figure 1: Human mobility data and their time series periodicity. (A) Human mobility data can be structured as algebraic tensors where each entry corresponds to the value of a specific variable $\gamma$ at a particular spatial location $n$ and time step $t$. The variable dimension $\gamma$ is flexible and can encode different aspects of mobility, such as inflow and outflow at a particular spatial location. Moreover, the temporal lags such as data from different years can also be incorporated as distinct variables within this tensor framework. (B) Illustration of weekly periodicity in passenger flow time series. This panel displays an example of hourly passenger flow time series that exhibits a clear weekly periodicity. This periodicity is visually highlighted by annotating the data points according to weekly cycles (e.g., blue and orange circles). The time series reveals distinct mobility patterns between weekdays and weekends, with weekdays showing more pronounced morning and afternoon peak hours. (C) Example time series of passenger flow for comparison. (D) Auto-correlation analysis through scatter plots. These panels present scatter plots that visualize the relationship between trip counts at a given time step $t$ and trip counts at previous time steps tkt-k, in which $k$ represents different time lags (e.g., $k=1,24,168$ correspond to hourly, daily, and weekly cycles, respectively). Points that align closely along the anti-diagonal in these scatter plots implies positive auto-correlation and potential periodicity. Notably, the scatter plot for a weekly cycle demonstrates a more remarkable similarity compared to the daily and hourly cycles, visually confirming the strong weekly periodicity in the example time series.
@@
Substantial progress has been made in understanding human mobility regularity and patterns from analytical perspectives such as entropy and predictability (Oliveira et al., 2016; Barbosa et al., 2018; Delussu et al., 2023). However, the existing methods are not well-suited to quantifying periodicity across spatiotemporal domains and discovering dynamic patterns that related to human mobility regularity. Recently, interpretable machine learning methods (Murdoch et al., 2019; Rudin et al., 2022; Brunton and Kutz, 2022) offer new opportunities to improve appearance, robustness, and interpretability of data-driven exploration in mobility data. In terms of model development, sparse linear regression (Tibshirani, 1996; Jenatton et al., 2011) can be used to reformulate the problem of sparse identification of nonlinear dynamical systems (Brunton et al., 2016), which laid the foundations of several machine learning frameworks that designed for variable and feature selection in scientific discovery. In that venue, sparsity is one of the most important aspects to reinforce interpretability of machine learning models (Murdoch et al., 2019; Rudin et al., 2022). However, existing sparse methods face critical limitations when applied to large-scale human mobility data. First, classical sparse regression methods are primarily designed for variable selection from univariate and multivariate data and struggle to scale to multidimensional data with spatial and temporal dependencies. Second, mobility data are inherently noisy, irregular, and heterogeneous across spatiotemporal domains, which challenges standard sparse models that lack built-in mechanisms for structured robustness. Third, even when sparsity is enforced, the resulting models do not yield interpretable or comparable metrics for periodicity strength across different spatial locations, variables, and time phases.
@@
To bridge these gaps, we propose a multidimensional sparse autoregression framework that models time series of mobility flows across space and time, explicitly captures dominant auto-correlations, and provides a unified and interpretable measure of periodicity strength at daily and weekly cycles. This interpretable machine learning framework builds upon the tensor data structure in the field of multi-linear algebra. Formally, human mobility data are spatially- and temporally-varying, and can be represented as a third-order tensor of non-negative integers $x_{n,\gamma,t}\in\{0,1,2,\cdots\}$ as shown in Fig. 1A, where $n\in[N]$ indexes spatial locations, $\gamma\in[\delta]$ indexes different variables or time phases, and $t\in[T_{\gamma}]$ indexes time steps. This algebraic tensor structure captures the multidimensional nature of mobility systems and aligns with a rich body of work on tensorial machine learning for spatiotemporal pattern discovery (Kolda and Bader, 2009; Chen et al., 2024). Each individual sequence $x_{n,\gamma,t}$ is a time series, and thus can be seamlessly characterized by using time series models such as autoregression (Hamilton, 2020). As exemplified by Fig. 1B, real-world passenger flow time series demonstrate weekly periodicity by picking data points with a weekly cycle. When these data points are denoted by $x_{t},t\in[T]$, the auto-correlations in Fig. 1D can be formulated as $x_{t}\approx w_{k}x_{t-k}$, measuring self-similarity in mobility over any time lag $k$. Observing Fig. 1D, data points of the example time series are very close to the anti-diagonal curve at a weekly cycle (i.e., $k=168$), implying that weekly periodicity is positively auto-correlated. Moreover, the “closeness” of data points to the anti-diagonal at index $k=168$ in Fig. 1D justifies the strength of weekly periodicity of time series, demonstrating that the time series in Fig. 1C is less periodic at a weekly cycle than Fig. 1B.
@@ 2 Machine Learning Formula
Figure 2: Illustration of time series autoregression and its multidimensional setting on human mobility data. (A) The $d$th-order autoregression on the univariate time series can be formulated as an optimization problem for finding the optimal coefficients. Here, the order $d$ is an integer, determining the number of auto-correlations. (B) Multidimensional time series autoregression. Note that the symbol $\times$” simply refers to the element-wise product between matrices. (C) The interpretability of time series autoregression can be reinforced by sparse and non-negative coefficients. The upper bound of the number of nonzero entries in the coefficient vector is set as an integer $\tau$, which is formally defined as the sparsity level. In mathematics, the cardinality of index set $\Omega$ (i.e., indices of nonzero coefficients) can be used to reformulate sparsity constraints. As a result, the decision variables in the constructed optimization problem include both coefficients and index set.
Interpretable machine learning provides a way to understand underlying patterns of complicated data and systems (Brunton et al., 2016; Murdoch et al., 2019). This study starts from the core idea of autoregression on time series, namely, the current value of the time series can be predicted as a linear combination of its past values (Hamilton, 2020). The objective is to find a set of coefficients that minimizes the sum of squared autoregressive errors (Fig. 2A) with predefined constraints such as sparsity and non-negativity on the coefficients as shown in Fig. 2C. Sparsity, enforced by the constraint $|\Omega|\leq\tau$ on the index set $\Omega$, limits the number of nonzero coefficients to a small integer $\tau$. This implies that only a few past time lags significantly contribute to predicting the current value. For instance, Fig. 1D shows that the trip counts at time $t$ are highly consistent with the trip counts at time t168t-168, revealing predictable and periodic patterns. By contrast, outliers on the scatter plots of Fig. 1D at the daily cycle, i.e., $k=24$, demonstrate less auto-correlated periodic patterns in the passenger flow time series as shown in Figs. 1B-C.
@@
where $w_{n,\gamma,k},n\in[N],\gamma\in[\delta],k\in[d]$ are the non-negative coefficients of this $d$th-order autoregression formula, allowing one to quantify the positive auto-correlations such as daily and weekly periodicity as the dominant ones in time series. The last constraint is the normalization on the non-negative coefficients, letting the sum of the coefficients be one. Since the optimization problem is equivalent to the $\ell_{0}$-norm induced sparse autoregression, it can be addressed by mixed-integer optimization solvers with exact solutions (Bertsimas et al., 2016, 2020, 2024; Tillmann et al., 2024). On such multidimensional tensor time series, we can discover the weekly periodicity of human mobility across different spatial locations and variables (or time phases).
@@ 3 Metro Passenger Flow Periodicity
To demonstrate the effectiveness of our framework in capturing periodicity from real-world mobility data, we begin with a case study of metro passenger flow data in Hangzhou, China. This dataset offers a clean and well-structured setting with strong known periodic patterns, making it an ideal testbed for validating the ability of the multidimensional sparse autoregression framework to recover meaningful temporal cycles—particularly weekly periodicity across different stations and flow directions. In addition, the optimized periodicity values of these passenger flow time series can be verified to be strictly consistent with auto-correlation analysis as shown in Fig. 1D.
@@
Figure 3: Identification of weekly periodicity in inflow and outflow time series from the Hangzhou metro passenger flow dataset (from January 2 to 25, 2019) with the multidimensional sparse autoregression method (sparsity level $\tau=4$). (A) Daily average inflow trips of 81 metro stations. Note that the trip records of the station #54 are missing. (B) Weekly periodicity of hourly inflow time series achieved by the sparse autoregression with sparsity level $\tau=4$. 64 stations show high periodicity ($\geq$ 0.8). The average periodicity value is 0.844. (C) Weekly periodicity of hourly outflow time series. 53 stations show high periodicity ($\geq$ 0.8). The average periodicity value is 0.829. Since inflow and outflow are two variables in the tensor data, their coefficients are optimized simultaneously with the same index set. (D) Scatter plots of trip counts at time $t$ versus time t168t-168 of 6 selected stations, illustrating varying strengths of weekly periodicity. The time series with a high weekly periodicity value also presents a strong similarity of data points between times $t$ and t168t-168. (E) Hourly inflow time series for stations #14 and #15, shown across three consecutive weeks: January 2–8, January 9–15, and January 16–22. Strong overlap among these weekly segments indicates high periodicity at a weekly cycle.
The metro passenger flow dataset in Hangzhou [^1] The anonymized trip records are available at https://doi.org/10.5281/zenodo.3145404. was processed as a tensor structure with 81 stations, 2 variables (i.e., inflow and outflow), and 576 hours (i.e., over 24 days). Weekly periodicity is quantified by the non-negative coefficients at index $k=168$ in the multidimensional sparse autoregression with the prescribed order $d=168$. In the experiments, we set the sparsity level as $\tau=4$ depending on how many dominant auto-correlations we expect to identify. In the mixed-integer optimization solver, the support set $\Omega$ was optimized simultaneously with the coefficients. In this case, the support set is optimized as $\Omega=\{1,24,144,168\}$, corresponding to hourly, daily, 6-day, and weekly cycles, respectively. The coefficients at index $k=168$ refer to as the strength of weekly periodicity. The analysis of weekly periodicity of metro passenger flow reveals the mobility patterns of both inflow and outflow variables. Table 1 summarizes the detailed comparison of weekly periodicity between inflow and outflow variables. Observing Figs. 3B-C and Table 1, the inflow time series at the end stations of metro lines (e.g., $\{27,28\}$, $\{30,31,32,34\}$, $\{35,36,37\}$, and $\{63,66,67\}$) showed stronger periodicity than the outflow time series. Conversely, the outflow time series of stations in downtown areas (e.g., $\{7,8,10,15,51\}$) were found to be more periodic than the inflow time series. These findings are relevant to commuting trips in urban areas.
@@
Table 1: Strength of weekly periodicity of the Hangzhou metro passenger flow data at the selected end stations of metro lines. The weekly periodicity of inflow data in most selected stations is stronger than outflow data.
Station # Inflow Outflow Station # Inflow Outflow Station # Inflow Outflow
23 0.77 0.65 31 0.90 0.70 39 0.73 0.74
24 0.46 0.48 32 0.91 0.74 40 0.83 0.88
25 0.24 0.54 33 0.90 0.80 62 0.95 0.82
26 0.78 0.62 34 0.90 0.76 63 0.89 0.77
27 0.93 0.78 35 0.91 0.76 64 0.89 0.81
28 0.87 0.69 36 0.88 0.74 65 0.86 0.80
29 0.77 0.66 37 0.85 0.74 66 0.87 0.73
30 0.81 0.71 38 0.76 0.75 67 0.93 0.79
@@
Fig. 3D shows the weekly periodicity values for six selected stations, measured as the strength of auto-correlations at a weekly lag (i.e., $k=168$) in the inflow time series. The source emphasizes that time series with high weekly periodicity values also show a strong similarity between data points at times $t$ and t168t-168, indicating positive auto-correlations. The weekly periodicity is visually confirmed by scatter plots and time series overlaps. For instance, station #14 shows a high weekly periodicity of 0.93, and its scatter plot shows a strong alignment along the anti-diagonal (Fig. 3D). In particular, the comparison between stations #14 and #15 in Fig. 3E answers why the strength of time series periodicity is a fair metric to reveal repeating patterns of human mobility. The periodicity value of station #15 is 0.71, which is smaller than station #14. As shown in Fig. 3E, the trip count time series at station #15 does not align perfectly across weeks, indicating lower weekly periodicity compared to the more consistent pattern observed at station #14. To summarize, on the passenger flow dataset, the comparable metric of weekly periodicity across different stations and inflow/outflow variables to discover periodic mobility patterns in a city-wide scale.